# Machine Learners Should Acknowledge the Legal Implications of Large Language Models as Personal Data

### v1

@@ 6 Summary
It is argued that LLMs do not contain personal data, and even if they did, extending data protection rights and obligations to LLMs would create unmanageable demands, particularly regarding rights to access and to be forgotten. Indeed, implementing these rights is challenging due to the nature and operational complexity of LLMs. Current research practices may be pushing the GDPR’s goal of technology neutrality to its limits.
However, practical difficulties in enforcing legal consequences do not negate the legal applicability. Reducing the scope of the GDPR based on unforeseen or conflicting interests would contradict its comprehensive protective purpose and intended technology neutrality. This could lead to significant legal uncertainty and potential liability for the ML community. Instead, solutions should focus on adapting legal consequences and developing new technical approaches for GDPR compliance.
If the ML community acknowledges the legal implications of LLMs as personal data, it can better engage with policymakers and other stakeholders to influence legislation. While a political shift in this direction is unlikely, the question of whether the GDPR needs to be adjusted for new AI technologies has already been raised (Sartor et al., 2020; Mitrou, 2018). Additionally, it remains to be seen how courts will address this issue (see pending cases such as (NOBY, 4)).
Overall, we encourage the ML community to acknowledge GDPR-related challenges in the development and deployment of LLMs. Joint efforts between computer science and law are needed to tackle the challenges.
@@ Acknowledgments
We thank Finn Fassbender and Thomas Klein for their helpful feedback on this paper.

### v2

@@ 6 Summary
It is argued that LLMs do not contain personal data, and even if they did, extending data protection rights and obligations to LLMs would create unmanageable demands, particularly regarding rights to access and to be forgotten. Indeed, implementing these rights is challenging due to the nature and operational complexity of LLMs. Current research practices may be pushing the GDPR’s goal of technology neutrality to its limits. However, practical difficulties in enforcing legal consequences do not negate the legal applicability. Reducing the scope of the GDPR based on unforeseen or conflicting interests would contradict its comprehensive protective purpose and intended technology neutrality. This could lead to significant legal uncertainty and potential liability for the ML community. Instead, solutions should focus on adapting legal consequences and developing new technical approaches for GDPR compliance.
If the ML community acknowledges the legal implications of LLMs as personal data, it can better engage with policymakers and other stakeholders to influence legislation. While a political shift in this direction is unlikely, the question of whether the GDPR needs to be adjusted for new AI technologies has already been raised [70, 55]. Additionally, it remains to be seen how courts will address this issue (see pending cases such as [58]). Finally, collaborative efforts from the machine learning community are necessary to tackle issues beyond memorization [43], in order to protect privacy, even though memorization continues to be a critical concern. Overall, we encourage the ML community to acknowledge GDPR-related challenges in the development and deployment of LLMs. Joint efforts between computer science and law are needed to tackle the challenges.
@@ Acknowledgments
We thank Finn Fassbender and Thomas Klein for their helpful feedback on this paper. Additionally, we would like to thank our anonymous reviewers at the Workshop on the Future of Machine Learning Data Practices and Repositories at ICLR 2025.